OpenStudy (anonymous):

The light turns green, and a bicyclist starts forward with an acceleration of 1.2m/s^2 . How far must she travel to reach a speed of 8.0m/s ?

OpenStudy (anonymous):

Quite simple... Use the 3rd Kinematic Equation \[v ^{2}=u ^{2}+2as\]Here; Initial velocity, u=0m/s Final velocity, v=8.0m/s Acceleration, a=1.2m/s^2 Displacement, s=? Substituting the values in the above equation; 8^2=0^2 + 2*1.2*s 64=2.4*s s=64/2.4 s=2.666m Therefore, the bicyclist must travel 2.666m to reach a speed of 8m/s.

OpenStudy (anonymous):

Thanks... Demitris for correcting me. I don't know about US or foreign education system but in Indian edu. system we can use any of those 3 equations and call them the 3 kinematic equation of motion. There is no such restriction on the usage of only first 2 equation.